OpenStudy (anonymous):

evaluate: (125)^−4/3

OpenStudy (anonymous):

125 = 5^3

OpenStudy (cruffo):

\[125^{-4/3}\] the negative power means flip and make the power positive. \[ = \frac{1}{125^{4/3}}\] The 3 in the denominator of the fraction is the root \[=\frac{1}{(\sqrt[3]{125})^{4}}\] Note that 5^3 = 125 \[=\frac{1}{5^{4}}\] \[=\frac{1}{625}\]